For the ff. write the half-reactions, over-all reaction and Nerst equation:

a. Microbial Fuel Cell
b. Alkaline battery
c. Ni-Cd battery
Answer #1

Microbial fuel cell.

Acetate fuel cell.

Anode reaction :

CH3COO- (aq) + 2H2O(l) - 8e   2CO2 (g) + 7H+ (aq)

Cathode reaction :

O2 (g) + 4 H+ (aq) + 4e 2 H2O (l)

Overall reaction

CH3COO- (aq) + H+ (aq)  + 2 O2 (g)   2CO2 (g) + 2 H2O (l) .

Nernst equation :

E = Eo - log

Acrivity of liquid is one.

Alkaline battery

Anode reaction

Zn (s) + 2 OH- (aq)   Zn(OH)2 (s) + 2e

Cathode reaction :

2MnO2 (s) + H2O (l) + 2e Mn2O3 (s) + 2OH- (aq)

Overall reaction :

Zn (s) + 2 MnO2 (s) + H2O(l)   Zn(OH)2 (s) + Mn2O3 (s) .

Nernst equation

E = Eo - log Q

Q is constant. As all reactants and products are either solid or liquid.

C.

Ni/Cd battery.

Anode reaction :

Cd (s) + 2OH- (aq) Cd(OH)2 (s) + 2e

Cathode reaction :

NiO(OH) (s) + H2O (l) + e   Ni(OH)2 (s) + OH- (aq)

Overall reaction :

Cd(s) + 2 NiO(OH) (s) + 2 H2O (l) Cd(OH)2 (s) + 2 Ni(OH)2 (s) .

Nernst equation :

E = Eo - log Q

Q is constant as reactant and product is either solid or liquid.